8) An object moved 200 m E, 50 m 30° N of W , then 20 m 45 ^ 0 of W. Find the final displacement


1
Expert's answer
2021-05-10T16:04:21-0400
"s=(200,0)+(-50\cos{30},50\sin{30})\\+(-20\cos{45},20\sin{45})=(142.6,39.1)\ m\\s=\sqrt{142.6^2+39.1^2}=148\ m\\
\theta=\arctan{\frac{39.1}{142.6}}=15\degree \text{ N of E}"